home *** CD-ROM | disk | FTP | other *** search
- Hubble Space Telescope, Daily and Status Reports
-
- TELNET STINFO.HQ.ESO.ORG or 134.171.8.4
- login: stinfo
-
- (extracted sample)
- HUBBLE SPACE TELESCOPE
- DAILY REPORT #900
-
- PERIOD COVERED: 0800 EST 06/02/93 - 0800 EST 06/03/93
- Daily Status Report as of 154/1200Z.
-
- 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED:
- 1.1 Completed FOC/96 4804 (UV Imaging of Nearby Galaxies with FOC/96)
-
- The Faint Object Camera (f/96) was used to make one ultraviolet
- image of nearby galaxy ESO-1348-3. The exposure was for ten minutes,
- and used filter F220W. The image was received and displayed. No
- anomalies were noted.
-
- 1.2 Completed GHRS/2 4065 (Dark Noise Monitoring GHRS)
-
- The High Resolution Spectrograph was used to make a series of 16
- Dark calibration images which will be used to monitor dark noise.
- All 16 observations were received and displayed in the OSS. They
- were reported to look okay.
-